OpenStudy (faithspring):

Solve to simplest form (2x^2 +25x +50) - (x^2 +2x +1)

OpenStudy (luckycoins888):

Final result : x2 + 23x + 49 Step by step solution : Step 1 : Equation at the end of step 1 : ((2x2 + 25x) + 50) - (x2 + 2x + 1) Step 2 : Trying to factor by splitting the middle term 2.1 Factoring x2+23x+49 The first term is, x2 its coefficient is 1 . The middle term is, +23x its coefficient is 23 . The last term, "the constant", is +49 Step-1 : Multiply the coefficient of the first term by the constant 1 • 49 = 49 Step-2 : Find two factors of 49 whose sum equals the coefficient of the middle term, which is 23 . -49 + -1 = -50 -7 + -7 = -14 -1 + -49 = -50 1 + 49 = 50 7 + 7 = 14 49 + 1 = 50 Observation : No two such factors can be found !! Conclusion : Trinomial can not be factored Final result : x2 + 23x + 49

OpenStudy (faithspring):

Thank you so much!